# Efficient first-principles inverse design of nanolasers

### v2

@@ 3 Computational Results
In this section, we demonstrate that our approach is efficient enough to attack fully three-dimension problems, since it is no more expensive than traditional linear inverse design. In particular, we consider an example nanolaser cavity designed for typical semiconductor-on-insulator platforms, in which a 220 nm semiconductor slab is patterned on top of a silica ($n=1.44$) substrate, with air above. The design region is otherwise similar to the 2d geometries, consisting of a 2d pattern and waveguide extruded with constant cross sections in the semiconductor layer, where voids in this layer are filled with air and the active material takes up the entire thickness (extrusion length) of the slab and is embedded within the semiconductor structure. PML+absorbing boundaries are used above and below the slab, separated from the semiconductor layer by $720$ nm. Additional implementation details are provided in App. D.
Using this 3d model, we again optimize nanolaser cavities with a gain medium size of $\sigma_{\text{g}}=500$ nm for the naive and nonlinear FOMs (Fig. 6), with no gain diffusion. Similar to the 2d case, the device optimized for the naive FOM creates a strong field concentration close to the cavity center, while the device optimized for the nonlinear FOM spreads the field more evenly over the active material region to counteract hole-burning effects. However, and in contrast to the 2d optimization results (Sec. 3.1), in 3d there is a more modest increase of the FOM when targeting nonlinear ($\text{FOM}/\zeta\approx 148$) vs the naive FOM ($\text{FOM}/\zeta\approx 90$) —there is an increase of a factor of $\approx 1.6$ (instead of the factor of 3 in Sec. 3.1). We attribute this result qualitatively to the fact that localizing a high-$Q$ resonance is more difficult in a 3d slab system than in 2d, because of the need to minimize out-of-plane radiation. That is, in 3d, more design freedom is devoted to increasing $Q$ than in 2d, leaving less design freedom to optimize the field profile for hole-burning effects. Nevertheless, the increase in the FOM is still a substantial improvement, with no additional computational cost over a naive approach. Post-processing the 3d results through an eigensolve again validates the high-$Q$ assumption for SPA-SALT and TCMT analysis, with values of $Q\approx 150$, while the optical extraction efficiency is moderately high $\geq 0.75$, lower than in 2d (Sec. 3.1) presumably due to out-of-plane radiation for this tiny ($\approx 2\lambda\times 2\lambda)$ cavity. We expect that higher $Q$ and efficiency should occur for larger design regions, similar to previous nanocavity designs.
@@ 4 Conclusion
While we believe that our new FOM should be immediately valuable to inverse design of nanolasers, it also offers a starting point for many future refinements. An obvious next step would be to replace our ad-hoc gain profile $D_{0}$ with a full first-principles model of the pumping process itself, since that process is likely to also depend somewhat on the geometry. For example, optical pumping can be modeled as a second Maxwell solve at the pump wavelength (e.g. for vertical illumination), with $D_{0}$ being computed from the pump absorption profile [4, 36, 65]. This is conceptually similar to the coupled pump–emission solves that have been applied to inverse design of Raman scattering [28, 29, 30], scintillation [31], and nonlinear harmonic generation [32, 33]. Moreover, one could include a geometry-dependent effect on the pump power to reach transparency (Sec. 2.1), as well as other effects such as absorption in the non-pumped region. Alternatively, modeling electrical pumping would involve solving a carrier-transport equation for $D_{0}$ [66], again coupled to the reciprocal lasing solve. Also, for the ultra-small active regions considered in nanolasers, pump blocking due to bandfilling effects become an important effect [10]. Even more complicated pumping models might involve nonlinear interactions between the pumping and emission processes, but since the emission should still simplify via a resonance approximation, we are hopeful that such effects will still be a tractable nonlinear solve (e.g. by a few Newton steps [39] for pre-selected poles), which would be differentiable by implicit adjoint methods [40]. In principle, optimization could also be applied to the full SALT equations, with the perturbative approach providing an initial guess that may reduce the difficulty of pole tracking—such a full-SALT optimization may become necessary in order if one wants to drive the system into the multi-mode regime. (Even without optimization, a SALT solve could be useful to quantify the accuracy of the SPA-SALT and TCMT approximations, although based on previous work we empirically expect the errors in such perturbative methods to be small for high-$Q$ single-mode cavities like the results in this paper.) Another effect that might be interesting to include in an optimization FOM is the lasing linewidth, for which accurate first-principles formulas in terms of the lasing mode have recently been derived that include the full effect of nano-patterned geometries [37], and which should be equally applicable to our reciprocal-solve estimate of the lasing mode in the high-$Q$ regime. It has already been experimentally demonstrated, that the proper distribution of active material in a nanostructured cavity can be used to significantly decrease the linewidth of a nanolaser [67].
@@
As we commented in Sec. 2.3.4, the “naive” FOM (15) that we employ for comparison purposes is actually a correct FOM for situations in which a single emitter is randomly distributed with probability density $\sim D_{0}$. This situation arises experimentally in systems employing quantum dots [68, 69], color centers [70, 71, 72], and similar isolated point-like emitters, whose placement is difficult to precisely control relative to the fabricated pattern. Despite its simplicity, Eq. (15) still encapsulates both the perturbative lasing physics and the out-coupling efficiency for such applications.
@@ Appendix A Comparison with SPA-SALT
In this appendix, we derive the connection between the single pole threshold approximation in Sec. 2.2 and SPA-SALT. In SPA-SALT notation, Let a passive cavity resonance $\Phi_{\mathrm{m}}(\mathbf{r})$ satisfy an equation similar to Eq. (4), but without the gain medium,
[(××${\left[(\nabla\times\nabla\times)-\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{\mathrm{m}}^{2}}{c^{2}}\right)\right]\Phi_{\mathrm{m}}(\mathbf{r})=0}.$ (A1)
Here, $\tilde{\omega}_{\mathrm{m}}$ is the complex frequency of the passive cavity resonance. Under the single-pole approximation, the lasing mode $\mathbf{E}_{L}(\mathbf{r})\approx a_{L}\Phi_{L}(\mathbf{r})$, where $a_{L}$ is the complex amplitude of the field. Thus, multiplying Eq. (A1) by $a_{L}$ and equating Eqs. (4, A1),
$\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{L}^{2}}{c^{2}}\right)\mathbf{E}_{L}(\mathbf{r})=\Big{(}\varepsilon_{c}(\mathbf{r})-i\Delta\varepsilon_{\Im}(\mathbf{r})\Big{)}\left(\frac{\omega_{L}^{2}}{c^{2}}\right)\mathbf{E}_{L}(\mathbf{r}).$ (A2)
@@
$\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{L}^{2}-\omega_{L}^{2}}{\omega_{L}^{2}}\right)\mathbf{E}_{L}(\mathbf{r})=-i\Delta\varepsilon_{\Im}(\mathbf{r})\mathbf{E}_{L}(\mathbf{r}),$ (A3)
@@
$\Delta\varepsilon_{\Im}(\mathbf{r})=\left(\frac{\gamma_{\perp}}{\gamma_{\perp}-i(\omega_{L}-\omega_{a})}\right)\frac{D_{0}(\mathbf{r})d}{1+\Gamma_{L}\left|\mathbf{E}_{L}(\mathbf{r})\right|^{2}},$ (A4)
@@
where as a reminder $\Gamma_{L}=\gamma_{\perp}^{2}/\left[\gamma_{\perp}^{2}+(\omega_{L}-\omega_{a})^{2}\right]$. Altogether, at the first lasing threshold, this yields
@@
$\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{L}^{2}-\omega_{L}^{2}}{\omega_{L}^{2}}\right)\mathbf{E}_{L}(\mathbf{r})=-i\left(\frac{\gamma_{\perp}}{\gamma_{\perp}-i(\omega_{L}-\omega_{a})}\right)d_{\textrm{thresh}}D_{0}(\mathbf{r})\mathbf{E}_{L}(\mathbf{r}).$ (A6)
@@
The SPA-SALT threshold prediction [2, 42] is then found by integrating both sides with respect to $\mathbf{E}_{L}(\mathbf{r})$ as SALT generally uses the biorthogonal normalization of the system’s lasing modes and resonances. We can instead recover Eq. (9) by restoring the assumption that $\omega_{L}=\omega_{a}$, $\tilde{\omega}_{L}=\omega_{L}-i\gamma_{0}$, and expanding $\frac{\tilde{\omega}_{L}^{2}-\omega_{L}^{2}}{\omega_{L}^{2}}$ to first order in $\gamma_{0}$, yielding
@@
$\varepsilon_{c}(\mathbf{r})\left(\frac{-2i\gamma_{0}}{\omega_{L}}\right)\mathbf{E}_{L}(\mathbf{r})=-id_{\textrm{thresh}}D_{0}(\mathbf{r})\mathbf{E}_{L}(\mathbf{r}),$ (A7)
which is then integrated by $\mathbf{E}_{L}^{*}(\mathbf{r})$.

### v3

@@ 3 Computational Results
In this section, we demonstrate that our approach is computationally efficient enough to attack fully three-dimension problems, since it is no more expensive than traditional linear inverse design. In particular, we consider an example nanolaser cavity designed for typical semiconductor-on-insulator platforms, in which a 220 nm semiconductor slab is patterned on top of a silica ($n=1.44$) substrate, with air above. The design region is otherwise similar to the 2d geometries, consisting of a 2d pattern and waveguide extruded with constant cross sections in the semiconductor layer, where voids in this layer are filled with air and the active material takes up the entire thickness (extrusion length) of the slab and is embedded within the semiconductor structure. PML+absorbing boundaries are used above and below the slab, separated from the semiconductor layer by $720$ nm. Additional implementation details are provided in App. E.
Using this 3d model, we again optimize nanolaser cavities with a gain medium size of $\sigma_{\text{g}}=500$ nm for the naive and nonlinear FOMs (Fig. 6), with no gain diffusion. Similar to the 2d case, the device optimized for the naive FOM creates a strong field concentration close to the cavity center, while the device optimized for the nonlinear FOM spreads the field more evenly over the active material region to counteract hole-burning effects. However, and in contrast to the 2d optimization results (Sec. 3.1), in 3d there is a more modest increase of the FOM when targeting nonlinear ($\text{FOM}/\zeta\approx 148$) vs the naive FOM ($\text{FOM}/\zeta\approx 90$) —there is an increase of a factor of $\approx 1.6$ (instead of the factor of 3 in Sec. 3.1). We attribute this result qualitatively to the fact that localizing a high-$Q$ resonance is more difficult in a 3d slab system than in 2d, because of the need to minimize out-of-plane radiation. That is, in 3d, more design freedom is devoted to increasing $Q$ than in 2d, leaving less design freedom to optimize the field profile for hole-burning effects. Nevertheless, the increase in the FOM is still a substantial improvement, with no additional computational cost over a naive approach. Post-processing the 3d results through an eigensolve again validates the high-$Q$ assumption for SPA-SALT and TCMT analysis, with values of $Q\approx 150$, while the optical extraction efficiency is moderately high $\geq 0.75$, lower than in 2d (Sec. 3.1) presumably due to out-of-plane radiation for this tiny ($\approx 1.5\lambda\times 1.5\lambda)$ cavity. We expect that higher $Q$ and efficiency should occur for larger design regions, similar to previous nanocavity designs.
@@ 4 Conclusion
While we believe that our new FOM should be immediately valuable to inverse design of nanolasers, it also offers a starting point for many future refinements. An obvious next step would be to replace our ad-hoc gain profile $D_{0}$ with a full first-principles model of the pumping process itself, since that process is likely to also depend somewhat on the geometry. For example, optical pumping can be modeled as a second Maxwell solve at the pump wavelength (e.g. for vertical illumination), with $D_{0}$ being computed from the pump absorption profile [csalt, I-SALT, siegman86]. This is conceptually similar to the coupled pump–emission solves that have been applied to inverse design of Raman scattering [ChristiansenMi20, PanCh21, YaoVe23], scintillation [Roques_Carmes_2022], and nonlinear harmonic generation [LinLi16, Mann2023]. Moreover, one could include a geometry-dependent effect on the pump power to reach transparency (Sec. 2.1), as well as other effects such as absorption in the non-pumped region. Alternatively, modeling electrical pumping would involve solving a carrier-transport equation for $D_{0}$ [chow1999semiconductor], again coupled to the reciprocal lasing solve. Also, for the ultra-small active regions considered in nanolasers, pump blocking due to bandfilling effects become an important effect [saldutti2023thresholdsemiconductornanolasers]. Even more complicated pumping models might involve nonlinear interactions between the pumping and emission processes, but since the emission should still simplify via a resonance approximation, we are hopeful that such effects will still be a tractable nonlinear solve (e.g. by a few Newton steps [EsterhazyLiu14] for pre-selected poles), which would be differentiable by implicit adjoint methods [Griewank2008]. In principle, optimization could also be applied to the full SALT equations, with the perturbative approach providing an initial guess that may reduce the difficulty of pole tracking—such a full-SALT optimization may become necessary if one wants to drive the system into the multi-mode regime. (Even without optimization, a SALT solve could be useful to quantify the accuracy of the SPA-SALT and TCMT approximations, although based on previous work and the numerical validation provided in App. D, we expect the errors in such perturbative methods to be small for high-$Q$ single-mode cavities like the results in this paper.) Another effect that might be interesting to include in an optimization FOM is the lasing linewidth, for which accurate first-principles formulas in terms of the lasing mode have recently been derived that include the full effect of nano-patterned geometries [pick], and which should be equally applicable to our reciprocal-solve estimate of the lasing mode in the high-$Q$ regime. It has already been experimentally demonstrated that the distribution of active material in a nanostructured cavity can be used to significantly decrease the linewidth of a nanolaser [fano].
@@
As we commented in Sec. 2.3.4, the “naive” FOM (15) that we employ for comparison purposes is actually a correct FOM for situations in which a single emitter is randomly distributed with probability density $\sim D_{0}$. This situation arises experimentally in systems employing quantum dots [pawel, QD], color centers [Lukin_color, Stas_2022, Sipahigil_2016], and similar isolated point-like emitters, whose placement is difficult to precisely control relative to the fabricated pattern. Despite its simplicity, Eq. (15) still encapsulates both the perturbative lasing physics and the out-coupling efficiency for such applications.
@@ Appendix A Comparison with SPA-SALT
A.1 Single mode lasing threshold
In this appendix, we derive the connection between the single pole threshold approximation in Sec. 2.2 and SPA-SALT. In SPA-SALT notation, Let a passive cavity resonance $\mathbf{e}_{1}(\mathbf{r})$ satisfy an equation similar to Eq. (4), but without the gain medium,
${\left[(\nabla\times\nabla\times)-\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{1}^{2}}{c^{2}}\right)\right]\mathbf{e}_{1}(\mathbf{r})}=0.$ (A1)
Here, $\tilde{\omega}_{1}$ is the complex frequency of the passive cavity resonance. Under the single-pole approximation, the lasing mode $\mathbf{E}_{L}(\mathbf{r})\approx a_{1}\mathbf{e}_{1}(\mathbf{r})/\sqrt{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{\mathrm{1}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}$, where $a_{1}$ is the complex amplitude of the field for the first lasing mode. Thus, multiplying Eq. (A1) by $a_{1}$ and equating Eqs. (4, A1),
$\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{1}^{2}}{c^{2}}\right)\mathbf{e}_{1}(\mathbf{r})=\Big(\varepsilon_{c}(\mathbf{r})-i\Delta\varepsilon_{\Im}(\mathbf{r})\Big)\left(\frac{\omega_{L}^{2}}{c^{2}}\right)\mathbf{e}_{1}(\mathbf{r}).$ (A2)
@@ A.1 Single mode lasing threshold
$\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{1}^{2}-\omega_{L}^{2}}{\omega_{L}^{2}}\right)\mathbf{e}_{1}(\mathbf{r})=-i\Delta\varepsilon_{\Im}(\mathbf{r})\mathbf{e}_{1}(\mathbf{r}),$ (A3)
@@
$\Delta\varepsilon_{\Im}(\mathbf{r})=\left(\frac{\gamma_{\perp}}{\gamma_{\perp}-i(\omega_{L}-\omega_{a})}\right)\frac{D_{0}(\mathbf{r})d}{1+\Gamma_{1}e_{c}^{-2}|a_{1}|^{2}\left|\mathbf{e}_{1}(\mathbf{r})\right|^{2}/\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{\mathrm{1}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega},$ (A4)
@@
where as a reminder $\Gamma_{1}=\gamma_{\perp}^{2}/\left[\gamma_{\perp}^{2}+(\omega_{L}-\omega_{a})^{2}\right]$. Altogether, at the first lasing threshold, this yields
@@
$\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{1}^{2}-\omega_{L}^{2}}{\omega_{L}^{2}}\right)\mathbf{e}_{1}(\mathbf{r})=-i\left(\frac{\gamma_{\perp}}{\gamma_{\perp}-i(\omega_{L}-\omega_{a})}\right)d_{\textrm{thresh}}D_{0}(\mathbf{r})\mathbf{e}_{1}(\mathbf{r}).$ (A6)
@@
The SPA-SALT threshold prediction [Ge_2010, cerjan_2016] is then found by integrating both sides with respect to $\mathbf{e}_{1}(\mathbf{r})$ as SALT generally uses the biorthogonal normalization of the system’s lasing modes and resonances. We can instead recover Eq. (9) by restoring the assumption that $\omega_{L}=\omega_{a}$, $\tilde{\omega}_{1}=\omega_{L}-i\gamma_{1}$, and expanding $\left(\tilde{\omega}_{L}^{2}-\omega_{L}^{2}\right)/\omega_{L}^{2}$ to first order in $\gamma_{1}$, yielding
@@
$\varepsilon_{c}(\mathbf{r})\left(\frac{-2i\gamma_{1}}{\omega_{L}}\right)\mathbf{e}_{1}(\mathbf{r})=-id_{\textrm{thresh}}D_{0}(\mathbf{r})\mathbf{e}_{1}(\mathbf{r}),$ (A7)
which is then integrated by $\mathbf{e}_{1}^{*}(\mathbf{r})$,
$d_{\textrm{thresh}}=\left(\frac{2\gamma_{1}}{\omega_{L}}\right)\left(\frac{\int_{\Omega}\mathbf{e}_{1}^{*}(\mathbf{r})\varepsilon_{c}(\mathbf{r})\mathbf{e}_{1}(\mathbf{r})d\Omega}{\int_{\Omega}\mathbf{e}_{1}^{*}(\mathbf{r})D_{0}(\mathbf{r})\mathbf{e}_{1}(\mathbf{r})d\Omega}\right).$ (A8)
A.2 Interacting threshold for a second lasing mode
To estimate the effect of spatial hole-burning on the thresholds of other potential lasing modes, we again approximate the spatial profile of each potential mode to be that of a different cavity resonance, $\mathbf{e}_{m}(\mathbf{r})$. We note that while the approximation of treating the first lasing mode as having the same spatial profile of a high-$Q$ resonance is quite good when the passive cavity resonance frequency is well-aligned with the central frequency of the gain (as we do here, see App. D), this same approximation on other lasing modes may be worse, especially if the resonance frequencies are outside of the gain window and thus experience significant line-pulling. Nevertheless, for the systems we consider here, which are tailored to support a single high-$Q$ passive cavity resonance, this approximation is still sufficient to estimate the range of single-mode behavior.
For any second lasing mode to reach threshold, such that the total lasing field can be decomposed into two different modes with distinct lasing frequencies $\mathbf{E}_{L}(\mathbf{r};t)\approx a_{1}\mathbf{e}_{1}(\mathbf{r})e^{-i\omega_{L}t}+a_{2}\mathbf{e}_{2}(\mathbf{r})e^{-i\omega_{L_{2}}t}$, we seek a non-trivial (i.e., for $|a_{2}|\ll 1$, but $|a_{2}|\neq 0$) solution to
$\varepsilon_{c}(\mathbf{r})\left(\frac{\tilde{\omega}_{2}^{2}-\omega_{L_{2}}^{2}}{\omega_{L_{2}}^{2}}\right)\mathbf{e}_{2}(\mathbf{r})=-i\left(\frac{\gamma_{\perp}}{\gamma_{\perp}-i(\omega_{L_{2}}-\omega_{a})}\right)\left(\frac{D_{0}(\mathbf{r})d^{(2)}_{\textrm{int thr}}}{1+\Gamma_{1}e_{c}^{-2}\left|a_{1}\right|^{2}\left|\mathbf{e}_{1}(\mathbf{r})\right|^{2}/\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{\mathrm{1}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\right)\mathbf{e}_{2}(\mathbf{r}).$ (A9)
Note that if $|a_{1}|=0$, the non-interacting threshold for this other mode is given by
$d^{(2)}_{\textrm{thr}}=i\left(\frac{\tilde{\omega}_{2}^{2}-\omega_{L_{2}}^{2}}{\omega_{L_{2}}^{2}}\right)\left(\frac{\gamma_{\perp}-i(\omega_{L_{2}}-\omega_{a})}{\gamma_{\perp}}\right)\left(\frac{\int_{\Omega}\mathbf{e}_{2}^{*}(\mathbf{r})\varepsilon_{c}(\mathbf{r})\mathbf{e}_{2}(\mathbf{r})d\Omega}{\int_{\Omega}\mathbf{e}_{2}^{*}(\mathbf{r})D_{0}(\mathbf{r})\mathbf{e}_{2}(\mathbf{r})d\Omega}\right),$ (A10)
thus
$d^{(2)}_{\textrm{int thr}}=d^{(2)}_{\textrm{thr}}\left[1+\Gamma_{1}e_{c}^{-2}|a_{1}|^{2}\frac{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{2}(\mathbf{r})|^{2}|\mathbf{e}_{1}(\mathbf{r})|^{2}d\Omega}{\left(\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{2}(\mathbf{r})|^{2}d\Omega\right)\left(\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}d\Omega\right)}\right].$ (A11)
To completely specify the interacting threshold of a second lasing mode within the SPA-SALT approximation (though noting that here we are using the Hermitian normalization and approximate orthogonality), we can solve for $|a_{1}|$ at arbitrary $d$ in the single mode regime up to and including the second interacting threshold using Eqs. A4 and A5 as
$\frac{d}{d_{\textrm{thresh}}}-1=\Gamma_{1}e_{c}^{-2}|a_{1}|^{2}\left[\frac{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}|\mathbf{e}_{1}(\mathbf{r})|^{2}d\Omega}{\left(\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}d\Omega\right)^{2}}\right].$ (A12)
To proceed, let
$\chi_{mn}=\frac{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{m}(\mathbf{r})|^{2}|\mathbf{e}_{n}(\mathbf{r})|^{2}d\Omega}{\left(\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{m}(\mathbf{r})|^{2}d\Omega\right)\left(\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{n}(\mathbf{r})|^{2}d\Omega\right)}.$ (A13)
Solving Eq. (A12) for $\Gamma_{1}e_{c}^{-2}|a_{1}|^{2}$ and into Eq. (A11), one can write an implicit equation for the interacting second modal threshold,
$d^{(2)}_{\textrm{int thr}}=d^{(2)}_{\textrm{thr}}\left(1+\left(\frac{d^{(2)}_{\textrm{int thr}}}{d_{\textrm{thresh}}}-1\right)\frac{\chi_{21}}{\chi_{11}}\right).$ (A14)
Finally, one can recover an explicit equation for the interacting thresholds of a second lasing mode as
$d^{(2)}_{\textrm{int thr}}=d^{(2)}_{\textrm{thr}}\left(\frac{1-\chi_{21}/\chi_{11}}{1-{d^{(2)}_{\textrm{thr}}\chi_{21}}/\left({d_{\textrm{thresh}}\chi_{11}}\right)}\right).$ (A15)
Note that this formula for the interacting threshold must be used with some caution, as without some care it can produce the misleading result of a negative interacting threshold. Instead, such a negative $d^{(2)}_{\textrm{int thr}}$ indicates that the mode never reaches threshold. To see, this, first note that by definition $d^{(2)}_{\textrm{thr}}>d_{\textrm{thresh}}$, and similarly, $\chi_{21}\leq\chi_{11}$, as the interacting modal overlap between two modes cannot be greater than the interacting overlap of a mode with itself. If $\chi_{21}=0$, i.e., the two modes do not spatially overlap, the interacting threshold of the second lasing mode is the same as its non-interacting value, $d^{(2)}_{\textrm{int thr}}=d^{(2)}_{\textrm{thr}}$. As $\chi_{21}$ is increased from zero, the ratio of $d^{(2)}_{\textrm{thr}}/d_{\textrm{thresh}}$ that yields a positive denominator becomes bounded. If, as a thought experiment, one could independently control $d^{(2)}_{\textrm{thr}}$ without altering $\chi_{21}$, increasing $d^{(2)}_{\textrm{thr}}$ above $d_{\textrm{thresh}}$ first yields a zero denominator, indicating a diverging $d^{(2)}_{\textrm{int thr}}$, before the interacting threshold becomes negative. Thus, within the SPA-SALT approximation,
$\frac{d^{(2)}_{\textrm{thr}}\chi_{21}}{d_{\textrm{thresh}}\chi_{11}}\geq 1\implies d^{(2)}_{\textrm{int thr}}\rightarrow\infty$. (A16)
In App. D we use the interacting thresholds to see compare the lasing thresholds of the first and second lasing modes.
A.3 Scalar residual for the SALT model
Using the expression in Eq. (A12) one can derive the lasing intensity for an arbitrary pumping strength above threshold $\Delta d=d-d_{\text{thresh}}$,
$|a_{1}|^{2}=\frac{\Delta d}{\Gamma_{1}e_{c}^{-2}\chi_{11}d_{\text{thresh}}}\,$, (A17)
which is the amplitude expression that we consider in App. B close to threshold ($\Delta d\ll 1$) and in the weak pumping regime ($|a_{1}|^{2}\ll 1$).
By using the expression for the lasing intensity derived in Eq. (A17) together with Eq. (3) in the single mode regime, it is possible to find the population inversion above threshold,
$D(\mathbf{r})=\left[1+\frac{d-d_{\text{thresh}}}{\chi_{11}d_{\text{thresh}}}\left|\mathbf{e}_{1}(\mathbf{r})\right|^{2}\right]^{-1}D_{0}(\mathbf{r})d\,$, (A18)
which can be directly inserted into Eq. (2). With this, we can define a normalized scalar residual measure based on the weak form of our finite element formulation that is a function of the pumping strength as
$r(d)=\frac{\int_{\Omega}|\nabla\times\mathbf{E}_{\text{r}}(\mathbf{r})|^{2}-\left[\varepsilon_{c}(\mathbf{r})-iD(\mathbf{r},d)\right]\left(\omega_{1}/c\right)^{2}\ |\mathbf{E}_{\text{r}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}{\int_{\Omega}\ \left(\omega_{1}/c\right)^{2}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\text{r}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\,$, (A19)
where we approximate the passive cavity mode with the reciprocal field in the high-$Q$ limit, while still considering the lasing frequency to be the real part of the cavity’s resonance frequency ($\omega_{L}\approx\omega_{1}$). Thus, the residual will account for errors in differences between the frequency of the mode and the one used to solve the reciprocal problem, as well as the error introduced through the nonlinear hole burning term. In App. D we evaluate this residual for different values of the pumping strength $d$ to validate the high-$Q$ approximation at threshold and above, which allows to see how spatial hole burning affects the designs optimized close to threshold ($\Delta d=d-d_{\text{thresh}}\ll 1$).